# Delayed Acceptance Markov Chain Monte Carlo for Robust Bayesian Analysis

### v2

@@ Abstract
This study introduces a computationally efficient algorithm, delayed acceptance Markov chain Monte Carlo (DA–MCMC), designed to improve posterior simulation in quasi-Bayesian inference. Quasi-Bayesian methods, which do not require fully specifying a probabilistic model, are often computationally expensive owing to the need to evaluate the inverse and determinant of large covariance matrices. DA–MCMC addresses this challenge by employing a two-stage process: In the first stage, proposals are screened using an approximate posterior, whereas a final acceptance or rejection decision is made in the second stage based on the exact target posterior. This reduces the need for costly matrix computations, thereby improving efficiency without sacrificing accuracy. We demonstrate the effectiveness of DA–MCMC through applications to both synthetic and real data. The results demonstrate that, although DA–MCMC slightly reduces the effective sample size per iteration compared with the standard MCMC, it achieves substantial improvement in terms of effective sample size per second, approximately doubling the efficiency. This makes DA–MCMC particularly useful for cases where posterior simulation is computationally intensive. Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis. …
@@ 1 Introduction
In a standard Bayesian inference, a probabilistic model is used for the data-generating process and the corresponding likelihood function, $p\left(\mathcal{D}|\boldsymbol{\theta}\right)$, where $\mathcal{D}$ denotes a set of observations with a sample size of $N$, $\boldsymbol{\theta}\in\mathbb{R}^{K}$ denotes a $K$-dimensional vector of unknown parameters, and $p\left(\boldsymbol{\theta}\right)$ denotes a prior. A posterior density is proportional to the product of the likelihood and prior, $p\left(\boldsymbol{\theta}|\mathcal{D}\right)\propto p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)$. Because a posterior is generally intractable, we simulate random draws from the density using numerical techniques, such as Markov chain Monte Carlo (MCMC) algorithms.
A well-known shortcoming of standard Bayesian analysis is that it needs to fully specify a probabilistic model. Alternative quasi-Bayesian approaches have been proposed to improve the robustness toward model misspecification, which have different constructions of quasi-likelihood that rely on loss function and moment conditions instead of a probabilistic model.
@@
The remainder of this paper is organized as follows: Section 2 introduces DA–MCMC, Section 3 compares the proposed algorithm with existing algorithm via synthetic and real data examples, and Section 4 concludes the paper.
@@ 2 Delayed acceptance MCMC
$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$
This structure improves computational efficiency by rejecting implausible proposals early, while still ensuring exact sampling from the full posterior distribution.
@@ 3.1 Synthetic data: Heteroskedastic linear regression
Model $N$ $K$ multiESS/iter

### v3

@@ Abstract
This study introduces a computationally efficient algorithm, delayed acceptance Markov chain Monte Carlo (DA–MCMC), designed to improve posterior simulation in quasi-Bayesian inference. Quasi-Bayesian methods, which do not require fully specifying a probabilistic model, are often computationally expensive owing to the need to evaluate the inverse and determinant of large covariance matrices. DA–MCMC addresses this challenge by employing a two-stage process: In the first stage, proposals are screened using an approximate posterior, whereas a final acceptance or rejection decision is made in the second stage based on the exact target posterior. This reduces the need for costly matrix computations, thereby improving efficiency without sacrificing accuracy. We demonstrate the effectiveness of DA–MCMC through applications to both synthetic and real data. The results demonstrate that, although DA–MCMC slightly reduces the effective sample size per iteration compared with the standard MCMC, it achieves substantial improvement in terms of effective sample size per second, approximately doubling the efficiency. This makes DA–MCMC particularly useful for cases where posterior simulation is computationally intensive. Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis.
@@ 1 Introduction
In standard Bayesian inference, a probabilistic model is used for the data-generating process and the corresponding likelihood function, $p\left(\mathcal{D}|\boldsymbol{\theta}\right)$, where $\mathcal{D}$ denotes a set of observations with a sample size of $N$, $\boldsymbol{\theta}\in\mathbb{R}^{K}$ denotes a $K$-dimensional vector of unknown parameters, and $p\left(\boldsymbol{\theta}\right)$ denotes a prior. A posterior density is proportional to the product of the likelihood and prior, $p\left(\boldsymbol{\theta}|\mathcal{D}\right)\propto p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)$. Because a posterior is generally intractable, we simulate random draws from the density using numerical techniques, such as Markov chain Monte Carlo (MCMC) algorithms.
A well-known shortcoming of standard Bayesian analysis is that it needs to fully specify a probabilistic model. Alternative quasi-Bayesian approaches have been proposed to improve the robustness toward model misspecification that rely on loss function and moment conditions instead of a probabilistic model.
@@
The remainder of this paper is organized as follows: Section 2 introduces DA–MCMC, Section 3 compares the proposed algorithm with an existing algorithm via synthetic and real data examples, and Section 4 concludes the paper.
@@ 2 Delayed acceptance MCMC
$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)+\left(1-r\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right)\delta_{\boldsymbol{\theta}^{\left(t\right)}}\left(\boldsymbol{\theta}^{\prime}\right),$
with
$r\left(\boldsymbol{\theta}^{\left(t\right)}\right)=\int\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)d\boldsymbol{\theta}^{\prime}.$
Here $\delta_{\boldsymbol{\theta}^{\left(t\right)}}\left(\cdot\right)$ denotes the Dirac mass at $\boldsymbol{\theta}^{\left(t\right)}$. This structure improves computational efficiency by rejecting implausible proposals early, while still ensuring exact sampling from the full posterior distribution.
@@ 3.1 Synthetic data: Heteroskedastic linear regression
Model NN $K$ multiESS/iter